I've come up with several solutions and not sure where I'm going wrong. Here is the question asked by my professor. My work is below but unsure of how to graph or what the actual plots are. Do I plot (2,3) or (2,0) and (0,3)
I appreciate your help
- Use either elimination or substitution to find the solutions (if any)
- Graph the two lines, labeling the x-intercepts, y-intercepts, and points of intersection
0.2x + 0.4y = 1.7 and 8.3x - 6.3y = -4.3
x/3 - y/5 = 4 and 3x/4 + 2y/3 = 6
4x = 8 and 5y = 15
4x = 8
5y = 15
4x/4 = 8/4
x = 2
5y = 15
5y/5 = 15/5
y = 3

Answer by Alan3354(69443)

- Use either elimination or substitution to find the solutions (if any)
- Graph the two lines, labeling the x-intercepts, y-intercepts, and points of intersection
0.2x + 0.4y = 1.7 and 8.3x - 6.3y = -4.3
x/3 - y/5 = 4 and 3x/4 + 2y/3 = 6
=======================
4x = 8 and 5y = 15
4x = 8
5y = 15
4x/4 = 8/4
x = 2
5y = 15
5y/5 = 15/5
y = 3
------
Plot the line x = 2. It's parallel to the y-axis and passes thru (2,0)
Plot the line y = 3. It's parallel to the x-axis and passes thru (0,3)
----
The intersection is the solution, (2,3)